OpenStudy (anonymous):

A computer can perform 30 identical tasks in six hours. At that rate, what is the minimum number of computes that should be assigned to complete 80 tasks within three hours?

OpenStudy (mbma526):

3 computers 1 computer 30 + 30=60 so add one more and you can 90 so 3

OpenStudy (mbma526):

you have 10 extra tasks

OpenStudy (anonymous):

unfortunately thats what I thought but that's incorrect =(

OpenStudy (mbma526):

really what would it be then?

OpenStudy (anonymous):

it said 6 but i dont know why

OpenStudy (mbma526):

hmm thats weird

OpenStudy (aric200):

30/6 = 5 tasks per hour 3*5 = 15 tasks in 3 hours. 15x = 80 =16/3 =5.33 Repeating The minimum number of computers would be 6 Computers.

OpenStudy (anonymous):

niceeeeeeeeeeee boss!!!!